# AoI-based Scheduling of Correlated Sources for Timely Inference

### v1

@@ Appendix A Proof of Theorem 1
We can further express the Bellman equation (A) for state $(\delta,1)$ at time $t$ as follows:
@@
Similarly, we can express the Bellman equation (A) for state $(1,\delta)$ as follows:
@@
Appendix B Computation of Penalty Function $f_{m}(\delta)$
Example 1.
$X_{m,t}$ $=a_{m}X_{m,t-1}+W_{m,t},$ (62)
[…]
where
$\bar{q}_{m,m}=q_{m,m}-\mathbf{q}_{m}Q^{-1}_{-1}\mathbf{q}_{m}^{T}$
$\mathbf{q}_{m}=[q_{m,1},\ldots,q_{m,m-1},q_{m,m+1},\ldots,q_{m,M}]$
Example 2.
$f_{m}(\delta)=\frac{\sum_{t=\delta+1}^{n}L(Y_{m,t},\phi(X_{m,t-\delta},Z_{-m,t-1}))}{n-\delta}.$ (64)
Appendix C Proof of Lemma 2
For the simplicity of presentation, we prove Lemma 2 for $M=3$ and $m=1$.
Part (a): we have
$g_{1}(\delta_{1},\delta_{2},\delta_{3})$
$=H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}})$
$\geq H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-1},X_{3,t-\delta_{3}})$
$\geq H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-1},X_{3,t-1})=f_{m}(\delta_{1}),$ (65)
where the inequalities hold due to [43, Lemma 12].
Part (b): We have
$g_{1}(\delta_{1},\delta_{2},\delta_{3})$
$=$ $H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}})$
$=$ $H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}})$
$+H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}})$
$\overset{(c)}{=}$ $H_{L}(Y_{1,t}|X_{1,t-\delta_{1}},X_{2,t-1},X_{3,t-1})$
$+I_{L}(Y_{1,t};X_{3,t-1}|X_{1,t-\delta_{1}},X_{2,t-1},X_{3,t-\delta_{3}})$
$+I_{L}(Y_{1,t};X_{2,t-1}|X_{1,t-\delta_{1}},X_{2,t-\delta_{2}},X_{3,t-\delta_{3}}),$
$\overset{(d)}{\leq}$ $f_{m}(\delta_{1})+\epsilon_{1,3}^{2}+\epsilon_{1,2}^{2}$
$\leq$ $f_{m}(\delta_{1})+2\max_{n\neq 1}\epsilon_{1,n}^{2}=f_{m}(\delta_{1})+O(\max_{n\neq 1}\epsilon_{1,n}^{2}),$ (66)
where (a) holds due to Markov chain $Y_{1,t}\leftrightarrow X_{2,t-1}\leftrightarrow X_{2,t-\delta_{2}}$ [43, Lemma 12], (b) is obtained by definition (9), (c) can be obtained by using similar steps of (a)&(b), and (d) holds because of (2). This completes the proof for $M=3$ and $m=1$. Following the same steps, we can prove it for any values of $M$ and $m$.
Appendix D Proof of Theorem 2
We define
Then, from (13)-(14) and (20)-(21), we have
@@ Appendix D Proof of Theorem 2
$\mathrm{V}_{opt}=$ $\inf_{\pi\in\Pi_{c}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{M}g_{m}(\Delta_{1}(t),\Delta_{2}(t),\ldots,\Delta_{M}(t))\right]$
$\overset{a}{=}$ $\inf_{\pi\in\Pi_{c}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))\right]$
$+\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{M}O\bigg{(}\max_{n=\neq m}\epsilon_{m,n}^{2}\bigg{)}$
$=$ $\mathrm{V}_{f,opt}+\frac{1}{1-\gamma}\sum_{m=1}^{M}O\bigg{(}\max_{n\neq m}\epsilon_{m,n}^{2}\bigg{)},$ (68)
where (a) is due to (18). Moreover, we can write
$\mathrm{V}_{\pi_{f,opt}}=$ $\mathbb{E}_{\pi_{f,opt}}\left[\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{M}g_{m}(\Delta_{1}(t),\Delta_{2}(t),\ldots,\Delta_{M}(t))\right]$
$\overset{b}{=}$ $\mathbb{E}_{\pi_{f,opt}}\left[\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{M}f_{m}(\Delta_{m}(t))\right]$
$+\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{M}O\bigg{(}\max_{n=\neq m}\epsilon_{m,n}^{2}\bigg{)}$
$=$ $\mathrm{V}_{f,opt}+\frac{1}{1-\gamma}\sum_{m=1}^{M}O\bigg{(}\max_{n\neq m}\epsilon_{m,n}^{2}\bigg{)},$ (69)
where (b) is due to (18).
By combining (D) and (D), we obtain (22).
Appendix E Proof of Theorem 3
By using triangle inequality, We can write
$\lim_{r\to\infty}|\mathrm{V}^{r}_{\mathrm{gain}}(\lambda^{*})-V^{r}_{opt}|$
$\leq$ $\lim_{r\to\infty}|\mathrm{V}^{r}_{\mathrm{gain}}(\lambda^{*})-\mathrm{V}^{r}_{f,\mathrm{gain}}(\lambda^{*})|+\lim_{r\to\infty}|\mathrm{V}^{r}_{f,\mathrm{gain}}(\lambda^{*})-V^{r}_{f,opt}|$
$+\lim_{r\to\infty}|V^{r}_{f,opt}-V^{r}_{opt}|.$ (70)
From (D), we obtain
$|V^{r}_{f,opt}-V^{r}_{opt}|=\frac{1}{1-\gamma}\sum_{m=1}^{rM}O\bigg{(}\max_{n\neq m}\epsilon_{m,n}^{2}\bigg{)}$ (71)
Similar to (D), we can also get
$\mathrm{V}^{r}_{\mathrm{gain}}(\lambda^{*})$
$=\mathbb{E}_{\pi_{gain}}\left[\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{rM}g_{m}(\Delta_{1}(t),\Delta_{2}(t),\ldots,\Delta_{M}(t))\right]$
${=}\mathbb{E}_{\pi_{gain}}\left[\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{rM}f_{m}(\Delta_{m}(t))\right]+\sum_{t=0}^{\infty}\gamma^{t}\sum_{m=1}^{rM}O\bigg{(}\max_{n=\neq m}\epsilon_{m,n}^{2}\bigg{)}$
$=\mathrm{V}^{r}_{f,gain}(\lambda^{*})+\frac{1}{1-\gamma}\sum_{m=1}^{rM}O\bigg{(}\max_{n\neq m}\epsilon_{m,n}^{2}\bigg{)}.$ (72)
If the global attractor condition holds, for gain index-based policy, from [37], we have
$\lim_{r\to\infty}|\mathrm{V}^{r}_{f,\mathrm{gain}}(\lambda^{*})-V^{r}_{f,opt}|=0.$ (73)
By substituting (71)-(73) into (E), we obtain Theorem 3.
Appendix F Proof of Theorem 4
This proves Theorem 4(a).
Part (b) To prove Theorem 4(b), the following Lemma is important.
Lemma 4.
For the penalty function $f_{m,k}$, we have
$\mathrm{Pr}\bigg{(}f_{m,k}(\delta)\in B_{m,k}(\delta)\bigg{)}=1-O(\eta).$ (74)
Lemma 5 (Hoeffding’s inequality).
Let $Z_{1},Z_{2},\ldots,Z_{n}$ are i.i.d. samples of a random variable $Z\in[0,1].$ For any $\bar{\eta}>0$, we must have
$\mathrm{Pr}\bigg{\{}\bigg{|}\mathbb{E}[Z]-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\bigg{|}\leq\sqrt{\frac{\mathrm{ln}(2/\bar{\eta})}{2n}}\bigg{\}}\geq 1-\bar{\eta}.$ (75)
Now, we will use Lemma 4, (VI-A)-(VI-A) to prove Theorem 4(b).
By using triangle inequality, we can have
$\bigg{|}J_{m,\lambda,{k}}(\delta)-J_{m,\lambda,{k-1}}(\delta)\bigg{|}$
$\leq|J_{m,\lambda,{k}}(\delta)-J_{m,\mathrm{optimistic},k}(\delta)|$
$~~~+|J_{m,\mathrm{optimistic},k}(\delta)-J_{m,\lambda,{k-1}}(\delta)|.$ (76)
By using [52, Theorem 12], we can have
$|J_{m,\mathrm{optimistic},k}(\delta)-J_{m,\lambda,{k-1}}(\delta)|$
$\overset{(a)}{\leq}\frac{1}{1-\gamma}\max_{\delta^{\prime}}d_{m,k-1}(\delta^{\prime}),$ (77)
where (a) holds due to Lemma 4 with probability $1-O(\eta)$.
Similarly, by using [52, Theorem 12], we have with probability $1-O(\eta)$:
$|J_{m,\lambda,{k}}(\delta)-J_{m,\mathrm{optimistic},k}(\delta)|$
$+\frac{1}{1-\gamma}\max_{\delta^{\prime}}d_{m,k-1}(\delta^{\prime})$ (78)
By substituting (F) and (F) into (F), we obtain Theorem 4(b).

### v2

@@ Appendix F Minimization of Time-averaged Sum of Inference Errors
We can further express the Bellman equation (F) for state $(\delta,1)$ at time $t$ as follows:
@@
Similarly, we can express the Bellman equation (F) for state $(1,\delta)$ as follows:
@@
Remark 2.
For larger values of $M$, we can have a gain index-based policy similar to discounted version of the problem, but with average cost version. To get the gain index of the average cost version, we determine the value function as follows:
The optimal relative value function for each source $m$ is given by
$h_{m,\lambda}(\delta)=f_{m}(\delta)-\mathrm{f_{m,opt,\lambda}}+\min\bigg\{\lambda,h_{m,\lambda}(\delta+1)\bigg\},$ (84)
where $\delta$ is the current AoI value and $h_{m,\boldsymbol{\lambda}}$ is the relative value function associated with the problem
@@
$\mathrm{f_{m,opt,\lambda}}\!\!=\inf_{\pi_{m}\in\Pi_{m}}\!\limsup_{T\to\infty}\!\frac{1}{T}\mathbb{E}\!\left[\sum_{t=0}^{T-1}f_{m}(\Delta_{m}(t))+\lambda\pi_{m}(t)\right],$ (85)
with $h_{m,\lambda}(1)=0$, $f(\delta)$ is defined in (14). The relative value function can be obtained by using relative value iteration algorithm [45]. The average cost gain index is given by
$\alpha_{m,\lambda}(\delta)=h_{m,\lambda}(\delta+1)-\lambda.$ (86)
For the average cost problem, the Lagrange update rule becomes
$\lambda_{k+1}=$ $\lambda_{k}+\frac{\theta}{k}\bigg(\frac{1}{T}\sum_{t=0}^{T-1}\sum_{m=1}^{M}\mathbf{1}\bigg(\alpha_{m,\lambda_{k}}(\Delta_{m}(t))>0\bigg)-N\bigg).$ (87)